OpenStudy (anonymous):

Help with the absolute convergence of infinite series cos(1/2n)/2^n

OpenStudy (anonymous):

\[\frac{ \cos \frac{ 1 }{2n} }{ 2^n }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes, but as an infinite series and if it converges conditionally or absolutely

OpenStudy (anonymous):

cos(1/2n) is always less than 1 so compare it to 1/2^n